### v1

# Monodromy Defects in Maximally Supersymmetric Yang-Mills Theories from Holography

Abstract
We study three Type II supergravity solutions which are holographically dual to codimension-2 supersymmetric defects in $(p+1)$-dimensional SU(NN) maximally supersymmetric Yang-Mills theories ($p=2,3,4$). In all of these cases, the defects have a non-trivial monodromy for the maximal abelian subgroup for the SO(9p9-p) R-symmetry. Such solutions are obtained by considering branes wrapping spindle configurations, changing the parameters (which alters the coordinate domain), and imposing suitable boundary conditions. We provide a prescription to compute the entanglement entropy of the effective theory on the defect. We find the resulting quantity to be proportional to the free energy of the ambient theory. A similar analysis is performed for the D5-brane wrapping a spindle, but we find that changing the coordinate domain does not lead to a defect solution, but rather to a circle compactification.
1 Introduction
The AdS/CFT correspondence [1, 2, 3] and its generalization to non-conformal theories [4, 5] have become a fruitful framework to study different aspects of supersymmetric gauge theories in various dimensions. Information about the dynamics of the gauge theory, operator spectrum and observables is geometrically encoded in the dual supergravity background, which allows for an exact or semi-analytic computation of these quantities.
Recently, there has been an increasing interest in the study of defects in quantum field theories. See for example [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] and references thereof. Of particular interest are defects in supersymmetric (conformal) theories which preserve some amount of superconformal symmetry, since this allows for tractable computations and to constraints for the observables.
Holographically, backgrounds dual to codimension-2 conformal monodromy defects preserving some amount of supersymmetry have been analysed in [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30] (in all these cases the ambient theory also preserves conformal symmetry). Such solutions are obtained at the level of lower dimensional gauged supergravities, and they share two important features: First, the metric asymptotes to $\text{AdS}_{d}$ written as a foliation of $\text{AdS}_{d-2}\times\text{S}^{1}$ over an interval. Second, the gauge fields have a non-trivial holonomy around the $\text{S}^{1}$. These are dual to background gauge fields for the abelian subgroup of the global symmetries, and the non-trivial holonomy of the supergravity gauge field leads to a non-trivial monodromy around the defect on the field theory.
Here, we focus on codimension-2 supersymmetric monodromy defects in (pp+1)-dimensional maximally supersymmetric SU($N$) SYM for [^1] The $p=3$ case has already been studied in [20, 24], but we review it here for completeness. $p=2,3,4$. We note that for $p\neq 3$ neither the ambient theory nor the defect preserve conformal symmetry. These solutions are obtained by considering the Dpp-branes wrapping spindles of [31], and changing the values of the integration constants, such that the coordinate domain of one of the coordinates of the spindle is now semi-infinite. This procedure was also applied in [30, 22, 23, 26, 24, 25].
A key element of our analysis is that in brane frame [32, 33], the near horizon metric of a Dpp-brane (for $p\neq 5$) is AdS. Although there is no conformal symmetry, as the Dilaton depends on the AdS radial coordinate, the fact that the metric in this frame is AdS, allows us to extend the same techniques used in the conformal cases mentioned above to the ones studied here. The case $p=5$ is analysed separately.
1.1 Brief Comments on Monodromy Defects
[…]
$\mathrm{d}s^{2}_{p+1}=\mathrm{d}x^{2}_{1,p-2}+\mathrm{d}\hat{\rho}^{2}+\hat{\rho}^{2}n^{2}\mathrm{d}z^{2},$ (1.1)
[…]
$\mathcal{A}_{R}=\mu_{R}\,\mathrm{d}z,\quad\mathcal{A}_{F_{\hat{I}}}=\mu_{\hat{I}}\,\mathrm{d}z,$ (1.2)
with $\hat{I}=1,...,\mathfrak{r}-1$. This is achieved by modifying the gauge covariant derivative $D$ as
$D_{\mu}\Psi\rightarrow\mathcal{D}_{\mu}\Psi=D_{\mu}\Psi-i\mathcal{R}[\Psi](\mathcal{A}_{R})_{\mu}\Psi-i\mathcal{F}_{\hat{I}}[\Psi](\mathcal{A}_{F_{\hat{I}}})_{\mu}\Psi,$ (1.3)
[…]
Since these gauge fields are singular at $\hat{\rho}=0$, the SO($1,p$) Lorentz symmetry is broken [^2] If both the ambient theory and the defect also preserve conformal symmetry, as is the case for $p=3$, then it is broken as $\rightarrow SO(2,p12,p-1)$ $\times$ SO(2). to SO(1,p21,p-2) $\times$ SO(2), as expected for a codimension-2 defect. The effect of this background gauge field is to make fields charged under these symmetries pick-up a non-trivial monodromy around the defect. Consider for example the gaugino ($\lambda$), which is only charged under the U(1)R. Under $z\rightarrow z+2\pi$, it will pick up an extra factor of $e^{-i\mathcal{R}[\lambda]}$.
[…]
$\mathrm{d}s^{2}_{p+1}=\hat{\rho}^{2}\left(\frac{\mathrm{d}x^{2}_{1,p-2}+\mathrm{d}\hat{\rho}^{2}}{\hat{\rho}^{2}}+n^{2}\mathrm{d}z^{2}\right).$ (1.4)
We then perform a Weyl transformation $\mathrm{d}s^{2}_{p+1}=\hat{\rho}^{2}\mathrm{d}\tilde{s}^{2}_{p+1}$. In the $p=3$ case, the SYM lagrangian does not change [^3] By this, we mean the SYM lagriangian in a generic curved space, which has the spacetime covariant derivative for the fermions and the $R\Phi^{2}$ coupling. This is the case since the Weyl rescaling (1.4) changes the background where the theory is formulated from a flat to a curved one., since the theory is conformally invariant. In the non-conformal cases, the Weyl transformation does not leave the SYM action invariant. The relevant term we are interested in is the Yang-Mills kinetic term, which after the Weyl rescaling reads
$S_{\text{YM}}=\int\mathrm{d}^{p+1}x\,\sqrt{-\tilde{g}}\,\left(-\frac{\hat{\rho}^{p-3}}{2g^{2}_{\text{YM}}}\tilde{g}^{\mu\nu}\tilde{g}^{\lambda\rho}\,\text{Tr}(F_{\mu\lambda}F_{\nu\rho})\right),$ (1.5)
so that the would be gauge coupling depends on the AdSp coordinate $\hat{\rho}$. We show below that this effective gauge coupling is matched by holography. We note that, in the conformal case, $\text{AdS}_{p}\times\text{S}^{1}$ precisely realises the spacetime symmetries of the defect, while in the non-conformal case, the factor of $\hat{\rho}^{p-3}$ break the $\text{AdS}_{p}$ isometries to SO(1,p21,p-2), matching the discussion above.
1.2 Organization of the paper
This paper is organized as follows: in Section 2 we start by reviewing general aspects of near horizon geometries to Dpp-branes and computing the central charge / free energy of the dual theories. We quickly review the brane frame of [32, 33] and find coordinates adapted to the study of defect solutions. In Section 3, we define the defect boundary conditions as in [20]. For $p=2,3,4$, we change the coordinate range of the spindle solutions in [34] to interpret them as dual to codimension-2 defects. Section 4 is dedicated to the study of defects on D5-branes, which do not enjoy an AdS metric in brane frame. Then, in Section 5 we provide a prescription to compute the defect entanglement entropy. This amounts to interpret the supergravity solutions as duals to a (p1p-1)-dimensional theory and applying the formulas of [35, 36]. To renormalize a divergence, we define a UV cut-off, for which the conformally AdS behaviour of the Dpp-brane ($p<5$) plays a central role. For $p=2,3,4$ we find the defect entanglement entropy to be proportional to the free energy of the ambient theory. This generalizes the results of [24] for conformal monodromy defects to non-conformal ones. We conclude in Section 6 with a summary of our results, and comments on lines of future research. An appendix with the supergravities used throughout this paper is provided.
2 Dpp-branes and the Dual Frame
In this section we review the Type II backgrounds holographically dual to (pp+1)-dimensional maximally supersymmetric SU($N$) Yang-Mills theories. These correspond to the decoupling limit of a stack of NN Dpp-branes. It is convenient to rewrite these solutions using the change of coordinates of [32, 33] that makes the metric conformally $\text{AdS}_{p+2}\times\text{S}^{8-p}$. It is this fact that allows us to generalize the study of codimension-2 monodromy defects in conformal theories to the maximally supersymmetric Yang-Mills theories in dimensions in which the theory is not conformally invariant. We also discuss a change of coordinates that matches the asymptotic behaviour of the monodromy defect solutions studied in the next section.
We start by considering the background configuration describing the backreaction of a single stack of NN Dpp-branes, which in string frame is given by [^4] For the D3-brane, this only gives the magnetic part of the self-dual 5-form. To solve the equations of motion, self-duality has to be imposed.
$\mathrm{d}s^{2}_{\text{st}}$ $=H_{p}(r)^{-\frac{1}{2}}\mathrm{d}x^{2}_{1,p}+H_{p}(r)^{\frac{1}{2}}\left(\mathrm{d}r^{2}+r^{2}\mathrm{d}s^{2}\left(\text{S}^{8-p}\right)\right),$ (2.1a)
$F_{8-p}$ $=(7-p)\ell^{7-p}\text{vol}\left(\text{S}^{8-p}\right),$ (2.1b)
$e^{\Phi}$ $=H_{p}(r)^{\frac{3-p}{4}},$ (2.1c)
where $\text{vol}(\text{S}^{8-p})$ is the volume form of the (8p)(8-p)-sphere, and in the decoupling limit
$H_{p}(r)=\frac{\ell^{7-p}}{r^{7-p}},\qquad\ell^{7-p}=(4\pi)^{\frac{5-p}{2}}\Gamma\left(\frac{7-p}{2}\right)N\,(\alpha^{\prime})^{\frac{7-p}{2}}g_{s}.$ (2.2)
There are $N$ units of Dpp-brane flux through the $\text{S}^{8-p}$
$\frac{1}{(2\pi\sqrt{\alpha^{\prime}})^{7-p}g_{s}}\int_{\text{S}^{8-p}}F_{8-p}=N.$ (2.3)
It was argued in [32, 33] that the metric (2.1a) is conformally $\text{AdS}_{p+2}\times\text{S}^{8-p}$. This can be seen explicitly using the change of coordinates
$\left(\frac{r}{\ell}\right)^{5-p}=\mathcal{R}^{2}\,(u\,\ell)^{2},\qquad\mathcal{R}^{2}=\left(\frac{2}{5-p}\right)^{2}.$ (2.4)
This change of coordinates exists for Dpp-branes for $p\leq 4$, we study the D5-brane case separately. In this coordinate the boundary is located at $u\rightarrow+\infty$ and the metric and dilaton are
$\mathrm{d}s^{2}_{\text{st}}$ $=\left(\mathcal{R}^{2}(u\,\ell)^{2}\right)^{-\frac{p-3}{2(p-5)}}\ell^{2}\left(\mathcal{R}^{2}\left(u^{2}\mathrm{d}x^{2}_{1,p}+\frac{\mathrm{d}u^{2}}{u^{2}}\right)+\mathrm{d}s^{2}\left(\text{S}^{8-p}\right)\right),$ (2.5a)
$e^{\Phi}$ $=\left(\mathcal{R}^{2}\,(u\,\ell)^{2}\right)^{\frac{(p-3)(p-7)}{4(p-5)}}.$ (2.5b)
In the Brane frame [^5] We change conventions with respect to [32, 33]., also called dual frame for non-conformal Dpp-branes
$\mathrm{d}s^{2}_{\text{dual}}=\left(\ell^{7-p}e^{\Phi}\right)^{-\frac{2}{7-p}}\mathrm{d}s^{2}_{\text{st}},\\$ (2.6)
it was argued [32, 5] that the coordinate $u$ is identified with the energy scale of the boundary theory. In fact, recalling that the effective (dimensionless) ’t Hooft coupling in $(p+1)$ dimensions is given in terms of the energy scale $E$ as
$\lambda_{\text{eff}}=g^{2}_{\text{YM}}\,NE^{p-3},\qquad g^{2}_{YM}=g_{s}(2\pi)^{p-2}(\alpha^{\prime})^{\frac{p-3}{2}}\\$ (2.7)
with $g^{2}_{\text{YM}}$ the Yang-Mills coupling constant, using the identification $u\sim E$, we can write a relation between the dilaton and the effective ’t Hooft coupling [^6] Here, the factor of $g_{s}$ on the right hand side is added manually when compared to [33], since $e^{\Phi_{\text{there}}}=g_{s}e^{\Phi_{\text{here}}}$.
$N\,g_{s}e^{\Phi}=c_{d}\lambda_{\text{eff}}^{\frac{7-p}{2(5-p)}},\\$ (2.8)
where $c_{d}$ can be found in equation (2.21) of [33]. In this frame it is also possible to read the boundary metric to be the flat $(p+1)$-dimensional Minkowski one. The gauge coupling can be determined from the DBI action, which we write using the dual frame metric
$S_{\text{DBI}}=\frac{\ell^{p-3}}{(2\pi)^{p-2}(\alpha^{\prime})^{\frac{p-3}{2}}g_{s}}\int\mathrm{d}^{p+1}x\,\sqrt{-\det(g_{\text{d,Ind}})}\,\left(e^{\Phi}\right)^{\frac{2(p-5)}{7-p}}\left(-\frac{1}{4}F_{\mu\lambda}F_{\nu\rho}(g_{\text{d,Ind}})^{\mu\nu}(g_{\text{d,Ind}})^{\lambda\rho}\right),$ (2.9)
where $g_{\text{d,Ind}}$ is the induced dual frame metric on the $(p+1)$-dimensional Minkowski space. From (2.6) and (2.5a)-(2.5b) we read $g_{\text{d,Ind}}=\mathcal{R}^{2}u^{2}\,\eta$, with $\eta$ the flat Minkowski metric. In (2.9), the gauge coupling is read by identifying the coefficient of $F^{2}$ with $1/(4g^{2}_{\text{YM}})$. In this case, using (2.8) one recovers (2.7), as was pointed out in [33].
The free energy [^7] We note that for even dimensional conformal theories the free energy matches the Weyl anomaly, which in holographic CFTs it is also equal to the central charge. of these theories can be computed using [35, 36] (this is computed using the string frame metric, see Section 5.1 for details)
$c_{\text{hol}}^{(p)}=\frac{2}{G_{N}}\frac{p^{p}\pi^{\frac{9-p}{2}}}{\Gamma\left(\frac{9-p}{2}\right)\left(9-p\right)^{p}}\left(\mathcal{R}^{2}(u\,\ell)^{2}\right)^{-\frac{(p-3)^{2}}{2(p-5)}}\ell^{8},\\$ (2.10)
where we have added the superscript $(p)$ as a label. We can write this expression in terms of the effective ’t Hooft coupling and $N$, using (2.5b) and (2.8). Omitting numerical factors, we find the free energy to be
$c_{\text{hol}}^{(p)}\sim N^{2}\lambda_{\text{eff}}^{\frac{p-3}{5-p}},\\$ (2.11)
which matches the supersymmetric localization computation [37] and its holographic version using spherical branes [38, 39]. We note that for Dpp-brane solutions, the statement above is true independently of the choice of parametrization for the holographic coordinate.
The advantage of using the parametrization in (2.5a)-(2.5b) where there is an explicit AdS factor, even if there is no manifest AdS isometry in the background, is that we can apply a similar analysis to those in [20, 21, 24] for codimension-2 monodromy defects in AdSd+2 geometries. The vacuum of these solutions corresponds to AdSd+2 written as a foliation of an interval over $\text{AdS}_{d}\times\text{S}^{1}$, and the defect solution is obtained as deformations of this vacuum. It is convenient for us to parametrize the Dpp-brane solutions above in this way and to discuss some general features of this coordinate choice, since the defect solutions studied in the following sections asymptote to the vacuum written in this form.
We can obtain such a parametrization of the pure Dpp-brane geometries, which correspond to the vacuum of the solutions studied in Section 3 below, that is, we move to coordinates adapted to the study of codimension-2 defect solutions by first writing the Minkowski metric as
$\mathrm{d}x^{2}_{1,p}=\mathrm{d}x^{2}_{1,p-2}+\mathrm{d}\tilde{y}^{2}+\tilde{y}^{2}\mathrm{d}z^{2},\\$ (2.12)
then, we perform a change of coordinates that mixes the holographic coordinate $u$ and the coordinate $\tilde{y}$ belonging to the field theory directions, $(u,\tilde{y})\rightarrow(\rho,y)$
$u=\rho y,\qquad\tilde{y}^{2}=\frac{1}{\rho^{2}y^{2}}(y^{2}-1),\\$ (2.13)
where $\rho>0$ and $y>1$. In these coordinates, for $p=2,3,4$ the AdS boundary (UV region) $u\rightarrow+\infty$ is mapped to $y\rightarrow+\infty$. The Dpp-backgrounds in dual frame now reads
$\mathrm{d}s^{2}_{\text{dual}}$ $=\mathcal{R}^{2}y^{2}\left(\rho^{2}\mathrm{d}x^{2}_{1,p-2}+\frac{\mathrm{d}\rho^{2}}{\rho^{2}}+\frac{1}{y^{2}(y^{2}-1)}\mathrm{d}y^{2}+\frac{y^{2}-1}{y^{2}}\mathrm{d}z^{2}\right)+\mathrm{d}s^{2}(\text{S}^{8-p}),$ (2.14a)
$F_{8-p}$ $=(7-p)\ell^{7-p}\text{vol}\left(\text{S}^{8-p}\right),$ (2.14b)
$e^{\Phi}$ $=\left(\mathcal{R}^{2}\,(y\,\rho\,\ell)^{2}\right)^{\frac{(p-7)(p-3)}{4(p-5)}},$ (2.14c)
Near $y=1$ the geometry shrinks smoothly [^8] We note that under $y=\sec(\theta)$, $\frac{1}{y^{2}(y^{2}-1)}\mathrm{d}y^{2}+\frac{y^{2}-1}{y^{2}}\mathrm{d}z^{2}=\mathrm{d}\theta^{2}+\sin^{2}(\theta)\mathrm{d}z^{2}.$ (2.15) Although the metric looks like a $\text{S}^{2}$, there is no SO(3) isometry since the warp factors and the dilaton depend on $\theta$.
$\frac{1}{y^{2}(y^{2}-1)}\mathrm{d}y^{2}+\frac{y^{2}-1}{y^{2}}\mathrm{d}z^{2}\underset{y\rightarrow 1}{\longrightarrow}\mathrm{d}y^{2}+y^{2}\mathrm{d}z^{2}.\\$ (2.16)
On the other hand, we can read the boundary metric at the UV region $y\rightarrow+\infty$, from the asymptotic expansion (here we use $\rho=1/\hat{\rho}$)
$\mathrm{d}s^{2}\rightarrow\frac{\mathrm{d}y^{2}}{y^{2}}+y^{2}\left(\frac{1}{\hat{\rho}^{2}}\left(\mathrm{d}x^{2}_{1,p-2}+\mathrm{d}\hat{\rho}^{2}\right)+\mathrm{d}z^{2}\right),\\$ (2.17)
so that the boundary metric is $\text{AdS}_{p}\times\text{S}^{1}$, which is conformally $(p+1)$-dimensional Minkowski
$\frac{1}{\hat{\rho}^{2}}\left(\mathrm{d}x^{2}_{1,p-2}+\mathrm{d}\hat{\rho}^{2}\right)+\mathrm{d}z^{2}=\frac{1}{\hat{\rho}^{2}}\left(\mathrm{d}x^{2}_{1,p-2}+\mathrm{d}\hat{\rho}^{2}+\hat{\rho}^{2}\mathrm{d}z^{2}\right).\\$ (2.18)
This exact behaviour is found in the conformal defect solutions of [20, 21, 24]. In that context, the holographic interpretation is that one takes a Weyl transformation of the flat Minkowski metric, so that the theory is now formulated in $\text{AdS}_{p}\times\text{S}^{1}$. In this case, the interpretation is similar, but since for $p=2,4$ the theories are non-conformal, the Weyl rescaling also induces a non-trivial profile for the gauge coupling, which can be read using the DBI action (2.9) to be [^9] This corresponds to the boundary value $y\rightarrow+\infty$. The leading coefficient in the asymptotic expansion corresponds to the gauge coupling [33]. $\hat{\rho}^{p-3}/g^{2}_{\text{YM}}$, which coincides with the field theory discussion around (1.5). We also note that the flat space axis located at $\hat{\rho}=0$ is mapped to infinity after the Weyl rescaling.
To obtain the background dual to a codimension-2 monodromy defect in maximally supersymmetric Yang-Mills in $(p+1)$ dimensions, we can extend the procedure of [30, 22, 23, 26, 24, 25] to non-conformal branes. This is, we start by considering the solutions from [34] of Dpp-branes ($p=2,4,5$) wrapped on spindles [^10] The $p=3$ case was studied in [24], but we briefly review it for consistency.. Then, by considering a different range of the spindle coordinates, we can interpret the solution as a codimension-2 monodromy defect. We discuss this in detail in Section 3 for $p=2,3,4$ and Section 4 for $p=5$.
3 Codimension-2 Monodromy Defects for $p=2,3,4$
We now proceed to the holographic study of codimension-2 monodromy defects in (pp+1)-dimensional maximally supersymmetric SU($N$) SYM. To obtain these solutions, we follow [30, 22, 23, 26, 24, 25], that is, we start by considering the Dpp-branes wrapped on spindle solutions of [34] and consider a different range of the spindle coordinates, which allows us to interpret the solutions as defects.
3.1 General Aspects of the Solutions and Boundary Conditions
We start by considering the solutions of Dpp-branes wrapped on a spindle of [34]. These are solutions of a U(1)r truncation of the $(p+2)$-dimensional maximal SO(9p9-p) gauged supergravity, here $\mathfrak{r}$ is the rank of SO(9p9-p), obtained by reducing the 10D theory [^11] For $p=2$, that is, the reduction on $\text{S}^{6}$, the resulting 4D maximally supersymmetric theory has ISO(7) as gauge group. Here we consider only the compact SO(7) subgroup. on $\text{S}^{8-p}$. The bosonic field content of these lower dimensional U(1)r theories are just the metric, $\mathfrak{r}$ U(1) gauge fields and a proper number of scalars depending on $p$. The metric of the solutions under consideration takes the form [^12] We perform a change of coordinates of the form $y\rightarrow y^{c}$ of the solutions of [34] and a redefinition of $H$ and $P$ such that all the backgrounds behave in the way presented here.
$\mathrm{d}s^{2}_{p+2}=F(\rho,y)\left(\rho^{2}\mathrm{d}x^{2}_{1,p-2}+\frac{\mathrm{d}\rho^{2}}{\rho^{2}}+\mathrm{d}s^{2}(\Sigma)\right),\\$ (3.1)
where
$\mathrm{d}s^{2}(\Sigma)=\frac{1}{P}\mathrm{d}y^{2}+\frac{P}{H}n^{2}\mathrm{d}z^{2},\\$ (3.2)
[…]
$A^{I}=a^{I}(y)\,\mathrm{d}z,\quad\Phi^{(p)}=\Phi^{(p)}(\rho,y),\\$ (3.3)
where $I=1,...,\mathfrak{r}$ and $\Phi^{(p)}$ denotes the sets of scalars of the ($p+2$)-dimensional theory. The explicit form of $F$, $H$, $P$, $a^{I}$ and the scalars depends explicitly on $p$. We note that the supersymmetric solutions of the U(1)r theories under consideration are characterized by $2\mathfrak{r}$ integration constants $(q_{I},\alpha_{I})$.
In the case of [34], the values of the integration constants were chosen such that the range of $y$ was taken between zeros of the function $P$, called them $y_{1}$ and $y_{2}$, that is $P(y)>0$ for $y_{1}<y<y_{2}$. Then, regularity conditions are imposed such that (3.2) corresponds to a spindle.
In order to interpret the solutions of [34] as defects, we follow the procedure of [30, 22, 23, 26, 24, 25]: we extend the range of $y$ to be semi-infinite, that is, $y\in[y_{\text{core}},+\infty)$, where $y_{\text{core}}$ is the greatest zero of the function $P$. In order to achieve this, we need to change the possible values of the integration constants. Furthermore, we need to impose defect boundary conditions as in [20, 21, 24, 25].
  • We allow the metric and scalars to be singular at $\rho=0$ and $\rho\rightarrow+\infty$, since those are present in the vacuum solutions (2.14a)-(2.14c).
  • At $y=y_{\text{core}}$, we require (3.2) to shrink smoothly. For this, besides $P(y_{\text{core}})=0$ we require $P^{\prime}(y_{\text{core}})\neq 0$ and $H(y_{\text{core}})\neq 0$. Then, regularity of the 2D space $\Sigma$ at $y=y_{\text{core}}$ leads to
    $\mathrm{d}s^{2}(\Sigma)\bigg|_{y=y_{\text{core}}}\rightarrow\mathrm{d}y^{2}+k^{2}y^{2}\mathrm{d}z^{2},\quad k^{2}=\frac{(P^{\prime}(y_{\text{core}}))^{2}}{4H(y_{\text{core}})}n^{2}.\\$ (3.4)
    Since $z\sim z+2\pi$, one must impose $k^{2}=1$. Also, we impose that the gauge fields $A^{I}$ vanish at the end of the space
    $A^{I}(y_{\text{core}})=0,\\$ (3.5)
    and scalars to be regular.
  • When $y\rightarrow+\infty$, the metric (in dual frame) asymptotes to the vacuum (2.18), with $dz^{2}\rightarrow n^{2}dz^{2}$. If $n\neq 1$ there is either a conical deficit or excess in the boundary theory. For the gauge fields we have
    $\frac{1}{2\pi}\int_{\text{S}^{1}}A^{I}\bigg|_{y\rightarrow+\infty}=\mu^{I}.\\$ (3.6)
    Here, $\mu^{I}$ are interpreted as background gauge fields for the U(1)r subgroup of the R-symmetry of the dual maximal theory. Importantly, $\mu^{I}dz$ cannot be gauged away. As in the conformal defects of [20, 21], only for $\mu^{I}\neq 0$ these solutions describe the insertion of a codimension-2 monodromy defect in the dual ($p+1$)-dimensional theory, located at $\hat{\rho}=0$ in flat space. However, in the cases studied here (for $p\neq 3$), neither the ($p+1$) theory nor the defect are conformally invariant.
    After the Weyl rescaling that sets the field theory metric to be $\text{AdS}_{p}\times\text{S}^{1}$, discussed below (2.18) for the vacuum, the background gauge fields $\mu^{I}dz$ have a non-trivial holonomy around the $\text{S}^{1}$. Also, similar to the conformal defects of [20, 21, 24], in the solutions studied below the conical singularity parameter $n$ is only related to the monodromy of the gauge field dual to the U(1) R-symmetry.
We now proceed to the study of the $(p+2)$-dimensional solutions of [34], imposing the defect boundary conditions. The procedure is systematic and it is as follows: first we impose (3.5) and (3.6), which allows to find an expression for the integration constants $q_{I}$ in terms of $y_{\text{core}}$ and $\mu^{I}$, that is
$A^{I}(y_{\text{core}})=0\quad\text{and}\quad\frac{1}{2\pi}\int_{\text{S}^{1}}A^{I}\bigg|_{y\rightarrow+\infty}=\mu^{I}\quad\Rightarrow\quad q_{I}=q_{I}(y_{\text{core}},\mu^{I}),\\$ (3.7)
then, using this expression for $q_{I}$ in the condition $P(y_{\text{core}})=0$, allows to obtain an expression of $y_{\text{core}}$ in terms of $\mu^{I}$
$P(y_{\text{core}})=0\quad\text{and}\quad q_{I}=q_{I}(y_{\text{core}},\mu^{I})\quad\Rightarrow\quad y_{\text{core}}=y_{\text{core}}(\mu^{I}).\\$ (3.8)
Finally, imposing $k^{2}=1$ in (3.4) leads to a constraint for the chemical potentials $\mu^{I}$.
3.2 4D Gauged Supergravity
The first configuration we study corresponds to the D2-brane wrapping spindle of [34], which is a solution of the 4D U(1)3 truncation of the maximal ISO(7) gauged supergravity [40]. Here, we slightly change conventions to match the form of the solutions in (3.1), such that the vacuum solution (in brane frame) takes the form (2.14a)-(2.14b). The solution takes the form
$\begin{split}\mathrm{d}s^{2}&=\frac{2^{14/3}}{3^{7/3}g^{7/3}}y^{1/3}H^{1/2}\rho^{1/3}\left(-\rho^{2}\mathrm{d}t^{2}+\frac{\mathrm{d}\rho^{2}}{\rho^{2}}+\frac{\mathrm{d}y^{2}}{P}+\frac{P}{H}n^{2}\mathrm{d}z^{2}\right),\\[5.69054pt] A^{I}&=\left(\alpha_{I}+\kappa\frac{4}{3g}\sqrt{1-\frac{\nu}{q_{I}}}\frac{q_{I}}{h_{I}}n\right)\mathrm{d}z,\\[5.69054pt] e^{\phi_{0}}&=\frac{2^{4/3}}{3^{2/3}g^{2/3}}y^{2/3}\rho^{2/3},\qquad e^{\phi_{I}}=\frac{2^{2/3}}{3^{1/3}g^{1/3}}\frac{\rho^{1/3}\sqrt{H}}{y^{1/3}h_{I}},\end{split}$ (3.9)
where
$P=H-y^{2}-\nu y^{2/3},\qquad H=h_{1}h_{2}h_{3},\qquad h_{I}=y^{4/3}+q_{I}.$ (3.10)
We have also included the BPS-parameter $\nu$ that completely breaks supersymmetry. Besides $\nu$, the configuration is characterized by six integration constants $(\alpha_{I},q_{I})$. Also, $n$ is a conical singularity parameter (recall $z\sim z+2\pi$), while $\kappa=\pm 1$. The vacuum solution is obtained by setting $(\alpha_{I},q_{I})=0$ and $n=1$. We provide the uplift of to Type IIA in Appendix A.5.
In what follows we focus on the supersymmetric defect solutions, so we set $\nu=0$. We also fix $\kappa=1$ for convenience, as it can always be set to this value by $z\rightarrow-z$.
As explained above, in order to interpret the solution as dual to a codimension-2 defect, we need to impose suitable boundary conditions. These were explained in 3.1. We start by imposing boundary conditions for the gauge field (3.7) and that $y_{\text{core}}$ is a zero of $P$ (3.8), from where
$q_{I}=-\frac{3g\mu^{I}y_{\text{core}}^{4/3}}{3g\mu^{I}+4n},\qquad y_{\text{core}}=\frac{\sqrt{3g\mu^{1}+4n}\sqrt{3g\mu^{2}+4n}\sqrt{3g\mu^{3}+4n}}{8n^{3/2}}.$ (3.11)
Then, the next step is to impose a smoothly shrinking geometry at $y=y_{\text{core}}$ (3.4). This, combined with (3.11) leads to a constraint for the background gauge fields $\mu_{I}$
$g\mu^{1}+g\mu^{2}+g\mu^{3}=2(n-1).$ (3.12)
Using (A.29), we can rewrite this constraint and (3.11) in terms of the background gauge field for the flavour symmetries $(\mu_{F_{1}},\mu_{F_{2}})$ and the R-symmetry $\mu_{R}$. First, we note that (3.12) only fixes the R-symmetry background gauge field
$\mu_{R}=\frac{2}{g}(n-1).$ (3.13)
From here, we see that a non-trivial background gauge field for the R-symmetry is only possible when there is a conical singularity in the dual field theory. This feature was already known for conformal monodromy defects in conformal theories [20, 21], and from (3.13) we see that it also applies to non-conformal defects in the maximally symmetric 3D theory, which is non-conformal.
Finally, we rewrite (3.11) as explained above. We find that it only depends on the flavour background gauge fields and the conical singularity parameter
$y_{\text{core}}=\frac{\sqrt{2(n+1)-g\mu_{F_{1}}-2g\mu_{F_{2}}}\sqrt{2(n+1)-g\mu_{F_{1}}+g\mu_{F_{2}}}\sqrt{2(n+1)+2g\mu_{F_{1}}+g\mu_{F_{2}}}}{8n^{3/2}}.$ (3.14)
We quickly present now some cases where the expressions get highly simplified.
$A^{1}=A^{2}=A^{3}$ Truncation
This sub-sector is obtained by fixing $q_{1}=q_{2}=q_{3}$. We note that in this case both flavour gauge fields $A_{F_{1}}=A^{1}-A^{2}$ and $A_{F_{2}}=A^{2}-A^{3}$ are set to zero, so that the only non-trivial gauge field is the R-symmetry one. Therefore, as seen above, there is a defect only if we allow for $n\neq 1$. In this case the solution is characterized by
$y_{\text{core}}=\frac{(1+n)^{3/2}}{2^{3/2}n^{3/2}},\qquad q_{1}=\frac{n^{2}-1}{4n^{2}},\qquad\alpha_{1}=\mu^{1}=\frac{2}{3g}(1-n).$ (3.15)
This sector preserves the same amount of supersymmetry as the general solution. We now comment on sub-sectors of enhanced supersymmetry.
Sub-sectors with Enhanced Supersymmetry
Setting some of the gauge fields to either be zero leads to more supersymmetry being preserved (at the 10D level). This can be understood simply by looking at the uplift in Appendix (A.5), as a trivial gauge fields restores a subgroup of the isometries of the internal manifold.
We focus on the simplest case $A^{2}=A^{3}=0$. This restores an $SO(4)=SU(2)\times SU(2)$ worth of isometries in the internal manifold, so we have four times as much supersymmetry. In this case, the parameters defining the solution read
$y_{\text{core}}=\frac{\sqrt{3-n}}{\sqrt{2}\sqrt{n}},\qquad q_{1}=\frac{3(n-1)}{2^{2/3}(3-n)^{1/3}n^{2/3}},\qquad\alpha_{1}=\mu^{1}=\frac{2}{g}(1-n).$ (3.16)
3.3 5D Gauged Supergravity
In this section we briefly review the 5D supergravity background dual to a codimension-2 defect on 4D $\mathcal{N}=4$ SU(NN) SYM. This solution was obtained in [41] and recently studied in [42, 43, 44] in the context of black hole near horizon geometries and spindle compactifications. The extension of this solution that allows to have a codimension-2 superconformal defect interpretation was investigated in [24], where this solution was deeply analysed. We limit ourselves to presenting the solution and reviewing the main features, since it completes our discussion for codimension-2 defects in Dpp-brane theories.
The background is a solution of the 5D U(1)3 gauged supergravity. The bosonic fields are the metric, three gauge fields $A^{I}$ ($I=1,2,3$) and two scalars $\phi_{1}$ and $\phi_{2}$. See Appendix A.4 for details. In the conventions of (3.1), supergravity background reads
[…]
Here, the $\text{AdS}_{3}$ factor is of unit radius metric and SO(2,2) is an exact isometry of the configuration. The functions $h_{I}$, $H$ and $P$ depend only on $y$ and are given by
$H=y^{-2}h_{1}h_{2}h_{3},\qquad P=H-y^{2}-\nu,\qquad h_{I}=y^{2}+q_{I}.\\$ (3.18)
As before, $\nu$ is a non-supersymmetric parameter and $\kappa=\pm 1$. Furthermore, using the uplift in Appendix A.4 and imposing flux quantization we find $g=1/\ell$, with $\ell$ given by $p=3$ in (2.2).
We now study defect boundary conditions for the supersymmetric solution, so we set $\nu=0$. In this case, the configuration preserved $\mathcal{N}=(0,2)$ supersymmetry [20]. Also, for convenience we take $\kappa=1$. Using (3.7) and (3.8) we find
$q_{I}=-\frac{g\,y^{2}_{\text{core}}\mu^{I}}{n+g\,\mu^{I}},\quad y_{\text{core}}=\frac{1}{n^{\frac{3}{2}}}(n+g\mu^{1})^{\frac{1}{2}}(n+g\mu^{2})^{\frac{1}{2}}(n+g\mu^{3})^{\frac{1}{2}}.$ (3.19)
Imposing now (3.4) on our expression of $y_{\text{core}}$ leads us to
$g\mu_{1}+g\mu_{2}+g\mu_{3}=(n-1),\quad\Rightarrow\quad\mu_{R}=\frac{1}{g}(n-1),$ (3.20)
where we have used the definition of the R-symmetry gauge field in (A.23). In terms of the flavour gauge field, we can write
$y_{\text{core}}=\frac{\sqrt{2(2n+1)-3g\mu_{F^{\prime}}}\sqrt{4(2n+1)+3g(\mu_{F^{\prime}}-2\mu_{F})}\sqrt{4(2n+1)+3g(2\mu_{F}+\mu_{F^{\prime}})}}{12\sqrt{6}n^{3/2}}.$ (3.21)
As in the previous 4D solution, this model admits some sub-sectors with enhanced supersymmetry [20]: in the case $A^{1}=A^{2}$ and $A^{3}=0$ the supersymmetry corresponds to ${\cal{N}}=(2,2)$. Also, setting $A^{1}=A^{2}=0$ the solution preserves Small ${\cal{N}}=(4,4)$.
3.4 6D Gauged Supergravity
We conclude this section by studying the solution dual to codimension-2 defects in the D4-brane theory. This solution, obtained at the level of the 6D U(1)2 gauged supergravity in [34], reads
[…]
As before $\kappa^{2}=1$, $n>0$, $y>0$ for the scalars to be well defined, and the positive functions $H$ and $P$ are given by
$H=y^{-4}h_{1}h_{2},\qquad P=H-y^{2}-\nu y^{-2},\qquad h_{I}=y^{4}+q_{I},\\$ (3.23)
where we included the non-BPS parameter $\nu$. Here, we have made a change of variables with respect to [34], such that setting $q_{I}=\alpha_{I}=\nu=0$, using the uplift in Appendix A.3 and moving to the dual frame, we recover the vacuum written in the form of (2.14a). Before proceeding to the study of defect boundary conditions, let us comment that this solution can be lifted to the 7D U(1)2 AdS gauged supergravity that has been deeply studied as spindle-compactification [45, 44] and defect [24, 22, 23]. Focusing on the defect interpretation, it was shown in [24, 22, 23] that it is possible define values of the integration constants such that the solution can be interpreted as a codimension-2 monodromy defect on a 6D dual CFT. Since this proves that it is possible to obtain a defect solution in the 6D supergravity theory, we briefly repeat the analysis in our parametrization for completeness. Below, we choose $\kappa=1$ for convenience, and we focus on the supersymmetric solution, so we set $\nu=0$.
We impose boundary conditions as explained in Section 3.1. From (3.7) we find
$q_{I}=-\frac{g\,y^{4}_{\text{core}}\mu^{I}}{2n+g\mu^{I}}.$ (3.24)
Then, (3.8) leads
$y_{\text{core}}=\frac{g}{2n}\sqrt{\mu^{1}+\frac{2n}{g}}\sqrt{\mu^{2}+\frac{2n}{g}}.$ (3.25)
from here we see that we must take $\mu^{I}>-2n/g$. Finally, setting $k^{2}=1$ in (3.4) we find
$\mu^{1}+\mu^{2}=-\frac{1}{g}(n-1).$ (3.26)
As before, we find it convenient to parametrize the solution in terms of the monodromies of the vector fields dual to the R-symmetry and flavour currents. Since $A_{\text{R}}=-A^{1}-A^{2}$ and $A_{\text{F}}=A^{1}-A^{2}$, we define $\mu_{\text{R}}=-\mu^{1}-\mu^{2}$ and $\mu_{\text{F}}=\mu^{1}-\mu^{2}$. In terms of these, we have
$y_{\text{core}}=\frac{1}{4n}\sqrt{(1+3n)^{2}-g^{2}\mu^{2}_{\text{F}}},\qquad\mu_{\text{R}}=\frac{1}{g}(n-1),$ (3.27)
that give us also the expression of $q_{I}$ substituting in (3.24). So, as in the previous cases, a non-zero background gauge field for the R-symmetry is only possible when there is a conical deficit in the dual field theory. This was also noted in [24] for the 7D uplift of the solution. Finally, using the uplift in Appendix A.3 we find $g=1/\ell$, with $\ell$ as in Section 2.
We now briefly illustrate two sub-sectors of the theory.
U(1) Truncation
One case in which we can find more explicit results is taking the flavour gauge field to vanish, which is achieved by $A^{1}=A^{2}$. In this case $\mu_{F}=0$ and
$y_{\text{core}}=\frac{1+3n}{4n},\qquad q_{1}=\frac{(n-1)(1+3n)^{3}}{2^{8}n^{4}}.$ (3.28)
Supersymmetry Enhancement Sector
In the case we set to zero one of the gauge fields, say $A^{2}=0$, the 10D the number of supercharges preserved is doubled. This was already observed at the level of the 7D solution [23] (recall that the 6D solution under consideration can be obtained by dimensional reduction from the 7D one). The main difference being that in 6D there are no AdS isometries, so there are not superconformal killing spinors. In this case, we have
$y=\frac{\sqrt{1+n}}{\sqrt{2}\sqrt{n}},\qquad q_{1}=\frac{n^{2}-1}{4n^{2}}.$ (3.29)
4 The $p=5$ case
As stated in Section 2, the D5-brane needs to be studied separately. This is due to the fact that the coordinate change (2.4) is not well defined in this case. In fact, the D5-brane metric is not conformally $\text{AdS}_{7}\times\text{S}^{3}$ but rather a linear dilaton background with metric conformal to $\mathbb{R}^{1,5}\times\mathbb{R}^{+}\times\text{S}^{3}$
$\mathrm{d}s^{2}_{\text{st}}$ $=r\,\ell\left(\frac{\mathrm{d}x^{2}_{1,5}}{\ell^{2}}+\frac{\mathrm{d}r^{2}}{r^{2}}+\mathrm{d}s^{2}(\text{S}^{3})\right),$ (4.1a)
$F_{3}$ $=2\ell^{2}\text{vol}(\text{S}^{3}),$ (4.1b)
$e^{\Phi}$ $=\frac{r}{\ell}.$ (4.1c)
Although the dual metric does not contain an AdS factor, which is what allowed us to extend the results of [24] to the $p=2,4$ branes, it is still possible extend the solution of [34] to have a semi-infinite range for one of the spindle coordinates (see Section 3.1).
We start our analysis differently: first, we consider the lift to 10D of the 7D gauged supergravity solution of [34], corresponding to a D5-brane wrapping a spindle. Then, as in the previous section, we change the range of the spindle coordinate $y$, which allows us to interpret the solution as a monodromy defect. We show below that the vacuum solution behaves differently when compared to the $p=2,3,4$ cases.
The solution of [34] is obtained in the 7D U(1)2 gauged supergravity. In this case, theory contains the metric, two U(1) gauge fields ($A^{I}$), and two scalars ($\lambda^{I}$), with $I=1,2$. The background configuration can be then written as
[…]
where $\kappa^{2}=1$ and $H$ and $P$ are given by
$H=h_{1}\,h_{2}\,,\qquad P=p^{2}\,H-y^{2}-\nu y\,,\qquad h_{I}=y+q_{I}\,$,\\ (4.3)
and they satisfy $H>0$ and $P>0$ for the metric to have the right signature, while the condition $p/h_{I}>0$ ensures that the scalars are real. As in the previous cases, we introduced the BPS-parameter $\nu$ that completely breaks supersymmetry. Setting $\nu=0$ leads to the supersymmetric solution of [34]. Also, $\alpha_{I}$ are constants related to the dual field theory background gauge field $\mu^{I}$.
At the level of the 7D solution, the solution is qualitatively different from the ones in Section 3, since it is not conformally $\text{AdS}_{5}\times\Sigma$, which is expected since, as explained above, the D5-brane vacuum does not have a $\text{AdS}_{7}$ factor in dual frame [^13] Recall that an $\text{AdS}_{7}$ can be written as a a foliation of $\text{AdS}_{5}\times\text{S}^{1}$ over an interval using (2.13)..
We focus our analysis on the supersymmetric solution, so from here we take $\nu=0$. We also set $\kappa=1$ and assume $n\geq 1$ for convenience. As explained in the previous sections, to obtain the defect solution we modify the range of the coordinate $y$ when compared to [34]. As in that case, it is convenient to use the symmetry.
$y\leftrightarrow-y,\quad q_{I}\leftrightarrow-q_{I},\quad p\leftrightarrow-p,\quad z\leftrightarrow-z,\quad$ (4.4)
to set $p>0$ and $h_{I}>0$. One difference with [34] appears when imposing the reality condition for $P(y)$. In that case, to obtain a bounded range of the coordinate $y$, it is necessary to impose $p<1$. Instead, to have a semi-infinite range for the coordinate $y$, we impose $p>1$. Finally, this symmetry also allows us to set $y_{\text{core}}\geq 0$, so that we have $y\in[y_{\text{core}},+\infty[$ are all positive values.
4.1 Vacuum Solution
Before studying defect boundary conditions, see Section 3.1, let us first analyse the vacuum of this configuration, obtained by setting $q_{1}=q_{2}=\alpha_{1}=\alpha_{2}=\nu=0$. The uplift to Type IIB Supergravity, see Appendix A.2, reads
$\mathrm{d}s^{2}_{\text{st}}$ $=p^{-\frac{5}{2}}e^{\frac{\rho}{2}}\sqrt{y}\left(\frac{p^{2}}{4g^{2}}\left(\mathrm{d}x^{2}_{1,3}+\mathrm{d}\rho^{2}+\frac{\mathrm{d}y^{2}}{y^{2}(p^{2}-1)}+(p^{2}-1)n^{2}\mathrm{d}z^{2}\right)+\frac{1}{g^{2}}\mathrm{d}s^{2}(\text{S}^{3})\right),$ (4.5a)
$F_{3}$ $=\frac{2}{g^{2}}\text{vol}(\text{S}^{3}),$ (4.5b)
$e^{\Phi}$ $=p^{-\frac{5}{2}}e^{\frac{\rho}{2}}\sqrt{y}.$ (4.5c)
comparing with (4.1b) we identify $g=1/\ell$. This condition do not change for the defect solution. Notice that, in brane frame, the $g_{zz}$ component does not depend on $y$, we comment on this feature below. The vacuum is parametrized in a similar way to (2.14a)-(2.14c), where the RG flow information is encoded both in $(y,\rho)$ via (2.13). In this case, it is also possible to move the vacuum to the canonical frame of via
$\rho=\frac{2}{p^{2}}\ln\left(\frac{r}{\ell}\right)-\sqrt{1-\frac{1}{p^{2}}}\tilde{y}+5\ln(p),\qquad y=\text{exp}\left(\sqrt{1-\frac{1}{p^{2}}}\tilde{y}+2\left(1-\frac{1}{p^{2}}\right)\ln\left(\frac{r}{\ell}\right)\right),$ (4.6)
which is the analogous of (2.13) for the $p=5$ case, however, the background does not take the form of (2.5a), that is, the canonical Dpp-brane solution with two of the spatial Minkowski directions written in polar coordinates, but rather
$\mathrm{d}s^{2}_{\text{st}}$ $=r\ell\left(\frac{p^{2}}{4}\left(\mathrm{d}x^{2}_{1,3}+\mathrm{d}\tilde{y}^{2}+(p^{2}-1)n^{2}\mathrm{d}z^{2}\right)+\frac{\mathrm{d}r^{2}}{r^{2}}+\mathrm{d}s^{2}(\text{S}^{3})\right),$ (4.7a)
$e^{\Phi}$ $=\frac{r}{\ell}.$ (4.7b)
Since in the solutions of our interest $z\sim z+2\pi$, the vacuum solution is dual to 6D maximally supersymmetric SU($N$) SYM on $\mathbb{R}^{1,4}\times\text{S}^{1}$. We note that this does not allow to interpretation of the extension of the solution of [34] as dual to a codimension-2 defect, since in order to do so, the vacuum should be of the form (2.14a), with the $(y,z)$ space being $\mathbb{R}^{2}$ in polar coordinates, but we instead have $\mathbb{R}\times\text{S}^{1}$. This changes the interpretation of $n$, in this case corresponds to the size of the $\text{S}^{1}$.
4.2 Boundary Conditions
We now proceed to the study of boundary conditions for solution. At $y=y_{\text{core}}$ these are the same as in 3.1, while for $y\rightarrow+\infty$, the condition for the gauge field remains the same, but the metric must asymptote to the vacuum studied in the section above.
As in the previous sections, we start by (3.7), from where
$q_{I}=-y_{\text{core}}\,(r_{I}+1),\quad r_{I}=-\frac{n\,p}{n\,p+2g\,\mu^{I}}$ (4.8)
Note that, since we are imposing $h_{I}>0$, which leads to $r_{I}y_{\text{core}}<0$. Since we are taking $y_{\text{core}}>0$, then $r_{I}<0$, which is equivalent to $\mu^{I}>-n\,p/(2g)$.
We now impose regularity for the metric at $y=y_{\text{core}}$. We start by demanding $P(y_{\text{core}})=0$. While in the previous cases this led to a relation between $y_{\text{core}}$ and the chemical potentials $\mu^{I}$, in this case we find
$P(y_{\text{core}})=y^{2}_{\text{core}}\left(p^{2}r_{1}r_{2}-1\right),\quad H(y_{\text{core}})=y^{2}_{\text{core}}r_{1}r_{2}.$ (4.9)
One possibility is to impose $y_{\text{core}}=0$. This however, also sets $H(y_{\text{core}})=0$, and since $P$ and $H$ go to zero with the same power of $y_{\text{core}}$, this leads to a finite size $\text{S}^{1}$ at $y=y_{\text{core}}$, so that near $y=y_{\text{core}}$ the $(y,z)$ space behaves as $\mathbb{R}\times\text{S}^{1}$ instead of $\mathbb{R}^{2}$, the later being the behaviour of our interest. In fact, one can check that setting $y_{\text{core}}=0$ leads to the vacuum solution. Instead, to find a defect solution we must fix
$p^{2}r_{1}r_{2}=1.$ (4.10)
In this way, the parameters characterizing the solution are $p$, $n$, $y_{\text{core}}$ and $r_{1}$, with $y_{\text{core}}>0$ and $r_{1}<0$. Of course, one could instead invert the relation between the parameters above to $(y_{\text{core}},r_{1})\rightarrow\mu^{I}$, to have the later ones as free parameters, but we find this unnecessary complicated.
We finish the regularity analysis by requiring smoothness condition for the metric at $y=y_{\text{core}}$, (3.4). Using (4.10), we find an expression for $r_{1}$ in terms of $p$ and $n$ provided
$p<\frac{1}{2}+\frac{1}{2}\sqrt{\frac{4+n}{n}},$ (4.11)
then, we find
$r_{1}=\frac{1}{n\,p^{3}}\left(-1-n\,p\pm\sqrt{1+2n\,p-n^{2}p^{2}(p^{2}-1)}\right).$ (4.12)
With these values, we can obtain a relation between the chemical potentials
$\mu^{1}+\mu^{2}=\frac{1}{g}.$ (4.13)
Interestingly, this combination of the chemical potentials does not depend on $n$, as was the case in the Dpp-branes with $p=2,3,4$.
With this, the background configuration asymptotes to a dual to 6D maximally supersymmetric SU($N$) SYM on $\mathbb{R}^{1,4}\times\text{S}^{1}$(4.5a)-(4.5c) and has a smoothly shrinking $\text{S}^{1}$ at $y=y_{\text{core}}$. Moreover, there are non-trivial background gauge fields for the U(1) R-symmetry and U(1) flavour symmetry. Hence, this solution can be interpreted describing the compactification of the 6D theory on $\text{S}^{1}$, with background gauge fields along the $\text{S}^{1}$. It would be interesting to see if this solution could be connected to the ones of [46, 47], in particular if one takes $A^{1}=A^{2}$.
5 Defect Entanglement Entropy
For backgrounds dual to codimension-2 conformal monodromy defects in conformal theories, it was showed in [24] that, by using a direct computation of entanglement entropy using [48], a divergence appears due to the non-compactness of the $y$ direction, for which the authors provide a renormalization procedure.
Here, we finish our study by applying the free energy formula [^14] This formula was proposed as a shortcut to the free energy computation via entanglement entropy. of [35, 36] to the codimension-2 monodromy defect solutions of Section 3. This requires choosing a holographic direction that realizes the holographic RG flow. For defect backgrounds studied here, which asymptote to the vacuum (2.14a)-(2.14c), the RG-flow information is encoded in the coordinates $(\rho,y)$, so that a direct computation of the central charge/ free energy of the $(p+1)$-dimensional theory with a defect is not possible in this parametrization.
By applying the formula of [35, 36] as if the supergravity backgrounds were duals to a (p1)(p-1)-dimensional theory, we find the same type of divergence as in [24]. We therefore extend their renormalization procedure to the cases studied here. A key element for this is that the backgrounds studied in [24] are of the form $\text{AdS}_{p}\times\text{S}^{1}$ foliated over an interval, with an asymptotic $\text{AdS}_{p+2}$ region at $y\rightarrow+\infty$. Although the solutions studied here, which are of the form (3.1)-(3.3), do not enjoy SO($2,p+1$) conformal symmetry (except for $p=3$), the metric in dual frame is of the exact same form as in [24]. It is this fact that allows us to extend the renormalization scheme to compute the entanglement entropy of the defect theory using backgrounds dual to codimension-2 non-conformal monodromy defects in non-conformal maximally supersymmetric SU($N$) theories.
With this, there is a clear procedure for the holographic computation of the codimension-2 defect free energy for $p=2,3,4$. We explain this now.
5.1 Renormalization Scheme
We start by reviewing the formula of [35, 36]. We consider backgrounds dual to a $(d+1)$-dimensional QFT, with string frame metric of the form
$\mathrm{d}s^{2}_{\text{st}}=a(\zeta,\theta^{i})\left(\mathrm{d}x^{2}_{1,d}+b(\zeta)\mathrm{d}\zeta^{2}\right)+g_{ij}(\zeta,\theta^{i})\mathrm{d}\theta^{i}\mathrm{d}\theta^{j}.$ (5.1)
Here, $\zeta$ is the holographic direction and $\theta^{i}$ are coordinates of the internal manifold. In terms of the following quantities
$V_{\text{int}}(\zeta)=\int\mathrm{d}\theta^{i}\sqrt{e^{-4\Phi}a(\zeta,\theta^{i})^{d}\det[g_{ij}]},\quad H=V_{\text{int}}^{2},$ (5.2)
with $\Phi$ the 10D dilaton, the central charge, or free energy for non-conformal theories, is defined as
$c_{\text{hol}}=\kappa_{d}d^{d}\frac{b(\zeta)^{\frac{d}{2}}H^{\frac{2d+1}{2}}}{G^{(10)}_{N}(H^{\prime})^{d}},$ (5.3)
[…]
We apply this formula for the solutions studied in Section 3, interpreting them solutions as duals to a (p1)(p-1)-dimensional QFT, hence treating $y$ and $z$ as part of the internal manifold. Using the uplifts in Appendix A, for $p=2,3,4$, we find
$V_{\text{int}}=(\mathcal{R}^{2})^{\frac{9-p}{2(5-p)}}\ell^{\frac{(7-p)^{2}}{5-p}}\rho^{\frac{p-1}{5-p}}\text{Vol}(\text{S}^{8-p})\int^{+\infty}_{y_{\text{core}}}\mathrm{d}y\,y^{\frac{p-1}{5-p}},$ (5.4)
which is clearly divergent. Here, we highlight that all the information about the defect, that is, the integration constants of the solutions, is entering only through $y_{\text{core}}$. Moreover, recall that for the vacuum solutions $y_{\text{core}}=1$.
[…]
As an example, let us comment on the UV regulator for the vacuum solutions (2.14a). In those backgrounds, the metric in dual frame is exactly AdSp+2 written as a foliation of $\text{AdS}_{p}\times\text{S}^{1}$ over an interval. We can use (2.13) to write the metric in the form of (2.5a). In this coordinate, the cut-off is simply defined as $u=\Lambda_{\text{UV}}$. Then, through (2.13), the regulator corresponds to $y=\Lambda_{\text{UV}}/\rho$ in the parametrization of (2.14a).
We now proceed to the definition of the UV cut-off for the defect solution. In brane frame these are of the form
$\mathrm{d}s^{2}_{p+2}=G(y)\left(\rho^{2}\mathrm{d}x^{2}_{1,p-2}+\frac{\mathrm{d}\rho^{2}}{\rho^{2}}+\frac{1}{P}\mathrm{d}y^{2}+\frac{P}{H}n^{2}\mathrm{d}z^{2}\right),\\$ (5.5)
where the functions $G$, $H$ and $P$ depend on the dimension but they are such that the metric is asymptotically $\text{AdS}_{p+2}$. The change of coordinate takes the metric to be asymptotically in FG form, which is the generalization of (2.13), was given in [48] and for (5.5) it takes the form
$u=\rho\,\text{exp}\left(\int\mathrm{d}y\,\sqrt{\frac{G-1}{P}}\right),\qquad\tilde{y}^{2}=\frac{1}{\rho^{2}}\,\text{exp}\left(2\int\mathrm{d}y\,\frac{1}{\sqrt{P(G-1)}}\right).$ (5.6)
In order to define the UV regulator, it is enough to consider an asymptotic expansion of this change of coordinates. We define $\mathcal{F}=\sqrt{\frac{G-1}{P}}$ and perform an asymptotic expansion as $y\rightarrow+\infty$, since this is the divergence we want to regulate [^16] Also, recall that in $(\rho,y)$ coordinates, the AdS boundary is at $y\rightarrow+\infty$. See the discussion in Section 2.. Up to next to leading order in powers of $y^{-1}$ we have
$\mathcal{F}=\frac{1}{y}+\frac{\mathcal{C}_{\alpha}}{y^{\alpha}}+\mathcal{O}\left(\frac{1}{y^{\alpha+1}}\right),$ (5.7)
with $\alpha>1$ and it is not necessarily an integer. Here we have used the fact that, for the solutions under consideration, as $y\rightarrow+\infty$ the functions $G\rightarrow y^{2}$, $H\rightarrow y^{4}$ and $P\rightarrow y^{4}$, so that the first order in the $\mathcal{F}$ expansion is always $y^{-1}$. With this, we can write the following asymptotic relation from (5.6)
$u\sim\rho\,y\,e^{\frac{\mathcal{C}_{\alpha}}{1-\alpha}y^{-\alpha+1}},\\$ (5.8)
which can be inverted using the Lambert $W$ function
$y\sim\mathcal{C}_{\alpha}^{\frac{1}{\alpha-1}}\,W\left(\mathcal{C}_{\alpha}\left(\frac{u}{\rho}\right)^{-(\alpha-1)}\right)^{-\frac{1}{\alpha-1}}.\\$ (5.9)
The UV cut-off can be then found by setting $u=\Lambda_{\text{UV}}$ and expanding for large $\Lambda_{\text{UV}}$. Up to next to leading order we define the regulator $y_{\text{UV}}$ as
$y_{\text{UV}}=\frac{\Lambda_{\text{UV}}}{\rho}\left(1+\frac{\mathcal{C}_{\alpha}}{\alpha-1}\left(\frac{\Lambda_{\text{UV}}}{\rho}\right)^{1-\alpha}\right).\\$ (5.10)
We highlight that this definition of the UV cut-off is universal for all metric of the form (5.5) that asymptotes to AdSp+2 as $y\rightarrow+\infty$, and not only to the solutions presented here. With this, we regulate the integral in (5.4) to obtain
$c^{\Lambda}_{\text{hol}}=\kappa_{d}\frac{2n}{G_{N}^{(10)}}\left(\frac{p-2}{p-1}\right)^{p-2}\frac{\pi^{\frac{11-p}{2}}(\mathcal{R}^{2})^{\frac{1}{2}}}{\Gamma\left(\frac{9-p}{2}\right)}\left(y_{\text{UV}}^{\frac{4}{5-p}}-y_{\text{core}}^{\frac{4}{5-p}}\right)\left(\mathcal{R}^{2}(\rho\ell)^{2}\right)^{-\frac{(p-3)^{2}}{2(p-5)}}\ell^{8},\\$ (5.11)
We note that the way in which this quantity depends on the energy scale $\rho\sim E$, that is, the factor $\left(\mathcal{R}^{2}(\rho\ell)^{2}\right)^{-\frac{(p-3)^{2}}{2(p-5)}}\ell^{8}$ also appears in the $(p+1)$-dimensional free energy (2.10) (up to $u\sim\rho$). This behaviour was already observed for the entanglement entropy of codimension-2 defect solutions, labeled “defect entanglement entropy” (dEE) in [24], where it was claimed that the dEE quantity is proportional to the central charge of the theory it is embedded in. Since the computation of [24] is contained within (5.11), we claim that for codimension-2 monodromy defects in $(p+1)$-dimensional maximally supersymmetric SU($N$) SYM, the dEE is proportional to the free energy of the $(p+1)$-dimensional theory. Following this observation we write
$c_{\text{hol}}^{\Lambda}=C_{p}\left(y_{\text{UV}}^{\frac{4}{5-p}}-y_{\text{core}}^{\frac{4}{5-p}}\right)c^{(p)}_{\text{hol}},\qquad C_{p}=\kappa_{d}\pi n\frac{(9-p)^{p}}{p^{p}}\frac{p-2}{p-1}{\cal{R}},\\$ (5.12)
where we denote by $c^{(p)}_{\text{hol}}$ the free energy/Weyl anomaly of the $(p+1)$-dimensional theory, which was obtained in (2.10).
Finally, in order to renormalize (5.11), we subtract the vacuum contribution. However, this has to be weighted with a factor of $n$, to take into account that defect solutions asymptote the vacuum (2.14a) with $dz^{2}\rightarrow n^{2}dz^{2}$ [24]. With this consideration, the renormalized quantity is given by
$c^{\text{renorm}}_{\text{hol}}=\lim_{\Lambda_{\text{UV}}\rightarrow+\infty}c^{\Lambda}_{\text{hol}}-n\,c^{\Lambda,\,\text{vac}}_{\text{hol}}.\\$ (5.13)
Since all the information about the defect/vacuum is contained in $y_{\text{UV}}$ and $y_{\text{core}}$, the relevant term in (5.13) is (recall that in the vacuum $y_{\text{UV}}=\Lambda_{\text{UV}}/\rho$ and $y_{\text{core}}=1$)
$c^{\text{renorm}}_{\text{hol}}\propto\lim_{\Lambda_{\text{UV}}\rightarrow+\infty}\left[y_{\text{UV}}^{\frac{4}{5-p}}-y_{\text{core}}^{\frac{4}{5-p}}-\left(\left(\frac{\Lambda_{\text{UV}}}{\rho}\right)^{\frac{4}{5-p}}-1\right)\right].\\$ (5.14)
For the UV part, using (5.10) and expanding for $\Lambda_{\text{UV}}\rightarrow+\infty$ we have
$y_{\text{UV}}^{\frac{4}{5-p}}$- $\left(\frac{\Lambda_{\text{UV}}}{\rho}\right)^{\frac{4}{5-p}}$ (5.15)
$=\left(\frac{\Lambda_{\text{UV}}}{\rho}\right)^{\frac{4}{5-p}}\left[1+\frac{4\mathcal{C}_{\alpha}}{(\alpha-1)(5-p)}\left(\frac{\Lambda_{\text{UV}}}{\rho}\right)^{1-\alpha}+\mathcal{O}\left(\left(\frac{\Lambda_{\text{UV}}}{\rho}\right)^{2-2\alpha}\right)\right]-\left(\frac{\Lambda_{\text{UV}}}{\rho}\right)^{\frac{4}{5-p}}.$
From here, we see that the leading order term is cancelled by the renormalization scheme. In order for the next to leading order term to be finite, it is required that
$\alpha\geq 1+\frac{4}{5-p}\\$ (5.16)
If the bound is saturated, there is a non-vanishing contribution to $c^{\text{renorm}}_{\text{hol}}$. We show below that, applying this procedure to the codimension-2 defects in Dpp-brane theories ($p=2,3,4$), in all the cases the bound is saturated. Finally, the dEE of the effective theory on a codimension-2 monodromy defect on $(p+1)$-dimensional maximally supersymmetric SU($N$) SYM reads
$c^{\text{renorm}}_{\text{hol}}=C_{p}\left(\mathcal{C}_{\alpha}\,\delta_{\alpha-1,\frac{4}{5-p}}-y_{\text{core}}^{\frac{4}{5-p}}+1\right)c_{\text{hol}}^{(p)}.\\$ (5.17)
We now proceed to the computation of (5.17) for the codimension-2 defects in Dpp-brane theories for $p=2,3,4$.
5.2 Supersymmetric Surface Defect in 4D SYM
We start with the computation of the dEE for the codimension-2 conformal defect in 4D $\mathcal{N}=4$ SU(NN) SYM, which was already computed in [24]. Nevertheless, we study this case for completeness, as it allows to support the results for the non-conformal defects.
In this case we have $G(y)=y^{\frac{2}{3}}H^{\frac{1}{3}}$, and together with (3.18) we find
$\alpha=3,\qquad\mathcal{C}_{3}=-\frac{q_{1}+q_{2}+q_{3}}{3}.\\$ (5.18)
[…]
$c^{\text{renorm}}_{\text{hol}}=C_{3}\left(-\frac{q_{1}+q_{2}+q_{3}}{3}-y_{\text{core}}^{2}+1\right)c_{\text{hol}}^{(3)}.\\$ (5.19)
Again, we write this expression in terms of the field theory observables, that is, the conical singularity parameter $n$ and the monodromies for the background gauge field of the flavour symmetries ($\mu_{F},\mu_{F^{\prime}}$)
$c^{\text{renorm}}_{\text{hol}}=\frac{\pi}{36n}\left(16(n-1)(5n+1)+3g^{2}\left(4\mu_{F}^{2}+3\mu_{F^{\prime}}^{2}\right)\right)c_{\text{hol}}^{(3)}.\\$ (5.20)
Interestingly, this quantity can be expressed in terms of observables intrinsic to the defect [24]: the defect Weyl anomaly $b$ and its conformal weight $h_{D}$. For backgrounds dual to supersymmetric codimension-2 conformal monodromy defects in 4D $\mathcal{N}=4$ SYM, these observables were computed in [20] using holographic renormalization. For our solutions
$h_{D}$ $=\frac{32(n-1)(1+2n)^{2}+27g^{3}\mu_{F^{\prime}}\left(\mu_{F^{\prime}}^{2}-4\mu_{F}^{2}\right)+18g^{2}(n+1)\left(4\mu_{F}^{2}+3\mu_{F^{\prime}}^{2}\right)}{2^{4}3^{3}n^{3}}c_{\text{hol}}^{(3)},$ (5.21)
$b$ $=\frac{\pi}{2^{3}3^{2}n^{2}}\left[32(n-1)\left(1+7n+19n^{2}\right)+27g^{3}\mu_{F^{\prime}}\left(\mu_{F^{\prime}}^{2}-4\mu_{F}^{2}\right)+18g^{2}(1+2n)\left(4\mu_{F}^{2}+3\mu_{F^{\prime}}^{2}\right)\right]c_{\text{hol}}^{(3)},$
from where it can be showed that the following relation holds
$c^{\text{renorm}}_{\text{hol}}=\frac{1}{3}\left(b-6\pi nh_{D}\right).\\$ (5.22)
Finally, we note that the sum of the $q_{I}$ is proportional to the conformal weight of the defect
$h_{D}=\frac{2}{3}\left(q_{1}+q_{2}+q_{3}\right)c_{\text{hol}}^{(3)}.$ (5.23)
5.3 Supersymmetric 3D Defect in 5D SYM
We now turn our attention to the defect solution in the D4-brane theory, obtained using the 6D gauged supergravity solution. In brane frame, we have $G(y)=y^{\frac{6}{5}}H^{\frac{1}{5}}$, and using (3.23) we obtain
$\alpha=5,\qquad\mathcal{C}_{5}=-\frac{2(q_{1}+q_{2})}{5}.\\$ (5.24)
Since in this case we have $p=4$ the bound (5.16) is saturated. The defect entanglement entropy (5.17) then reads
$c^{\text{renorm}}_{\text{hol}}=C_{4}\left(-\frac{2(q_{1}+q_{2})}{5}-y_{\text{core}}^{4}+1\right)c_{\text{hol}}^{(4)},\\$ (5.25)
which, in terms of the field theory parameters ($n,\mu_{F}$), is given by
$c^{\text{renorm}}_{\text{hol}}=\frac{C_{4}}{2^{8}5n^{4}}\Big((n-1)\left(1+n\left(29+n\left(227+767n\right)\right)\right)+2(1+3n)(1+11n)g^{2}\mu_{F}^{2}-g^{4}\mu_{F}^{4}\Big)c_{\text{hol}}^{(4)}.$ (5.26)
Maximally supersymmetric Yang-Mills in dimensions different from four fails to be a conformal theory due to the gauge coupling being a dimensionful parameter. However, as explained in [33], if one allows the gauge coupling to transform under conformal/Weyl transformations, as if it were a spurion, one finds a generalized conformal structure. Therefore, it is natural to wonder if the defect entanglement entropy (5.26) could be decomposed in terms of quantities intrinsic to the defect as in the 4D (conformal) case, that is, if it can be written in terms of the would be conformal weight of the defect under the generalized conformal structure and the defect free energy.
In order to achieve the decomposition described above, it is convenient to recall that 5D maximally supersymmetric Yang-Mills corresponds to the 6D $\mathcal{N}=(2,0)$ superconformal theory reduced on a circle. The holographic dual of the codimension-2 monodromy defect on the 6D $\mathcal{N}=(2,0)$ theory was studied in [24], where it was showed that the defect entanglement entropy is decomposed is the same fashion as (5.22). Then, a circle reduction of those quantities, using the techniques of [49], can be used to map the 6D quantities to 5D ones. We start by noting that (5.26) exhibits the same dependence on the defect parameters $(n,\mu_{F})$ as the analogous quantity given in eq. (5.2) of [24]. We elaborate more on this in the conclusion,
5.4 Supersymmetric Line Defect in 3D SYM
We end by studying the codimension-2 defect in the D2-brane theory. In this case, by moving to brane frame we find $G(y)=y^{\frac{6}{7}}H^{\frac{2}{7}}$. Then, from (3.10) we find
$\alpha=\frac{7}{3},\quad\mathcal{C}_{\frac{7}{3}}=-\frac{5}{14}(q_{1}+q_{2}+q_{3}).$ (5.27)
Interestingly, although the next to leading order term in (5.7) is not an integer power of $1/y$, the bound (5.16) is saturated. As stated before, this leads to an extra term in the defect entanglement entropy (5.17), which reads
$c^{\text{renorm}}_{\text{hol}}=C_{2}\left(1-\frac{5}{14}(q_{1}+q_{2}+q_{3})-y_{\text{core}}^{\frac{4}{3}}\right)c_{\text{hol}}^{(2)}.$ (5.28)
We now write this quantity in terms of the field theory parameters: the background gauge fields of the flavour $(\mu_{F_{1}},\mu_{F_{2}})$ and R-symmetry $\mu_{R}$. For this we use (3.11)-(3.14), which lead to
$c^{\text{renorm}}_{\text{hol}}$ $=C_{2}\frac{1}{224\Delta_{1}^{1/3}\Delta_{2}^{1/3}\Delta_{3}^{1/3}n^{2}}\bigg((g\mu_{F_{1}}-g\mu_{F_{2}}-2)(2g\mu_{F_{1}}+g\mu_{F_{2}}+2)(g\mu_{F_{1}}+2g\mu_{F_{2}}-2)-232n^{3}$ (5.29)
$-456n^{2}+54n\left(g\mu_{F_{1}}^{2}+g\mu_{F_{1}}g\mu_{F_{2}}+g\mu_{F_{2}}^{2}-4\right)+224n^{2}\Delta_{1}^{1/3}\Delta_{2}^{1/3}\Delta_{3}^{1/3}\bigg)c_{\text{hol}}^{(2)},$
where we used the auxiliary variables
$\Delta_{1}=2(1+n)-g\mu_{F_{1}}-2g\mu_{F_{2}},\qquad\Delta_{2}=2(1+n)-g\mu_{F_{1}}+g\mu_{F_{2}},\qquad\Delta_{3}=2(1+n)+g\mu_{F_{1}}+2\mu_{F_{2}}.$ (5.30)
6 Conclusion and Future Directions
In this paper, we computed the defect entanglement entropy for codimension-2 monodromy defect in maximally supersymmetric SU($N$) Yang-Mills theories in $(p+1)$-dimensions, ($p=2,3,4$). For $p\neq 3$ neither the ambient theories nor the defects are conformally invariant, thus generalizing the results of [24].
The defect geometries are obtained by considering the solutions of Dpp-brane wrapping a spindle of [34] and changing the range of one of the spindle coordinates, such that now it takes values on a semi-infinite interval. Defect boundary conditions are imposed, so that at one end the geometry shrinks smoothly, while asymptotically the metric goes to the Dpp-brane one, but we allow the gauge fields to be asymptotically constant. This constant is related to the monodromy of background gauge fields in the dual field theory.
In order to compute the defect entanglement entropy, we use the formula of [35, 36], interpreting the backgrounds as duals to a (p1)(p-1)-dimensional theory. This leads to a divergence which needs to be renormalized. To do so, we take advantage of the fact that in brane frame the metric asymptotes to $\text{AdS}_{p+2}$ written as a foliation of $\text{AdS}_{p}\times\text{S}^{1}$ over an interval. We provide a renormalization prescription that allows to obtain a finite result, similar to the one of [24]. This amounts to define a UV cut-off and subtract the contribution of the vacuum. This procedure is only possible for $p=2,3,4$, since the D5-brane does not have a $\text{AdS}_{p+2}$ asymptotic region in brane frame [33].
In the conformal case ($p=3$), the renormalization procedure leads to a quantity that is proportional to a linear combination of the Weyl anomaly on the defect and its conformal weight. This feature is common to other conformal monodromy defects as showed in [24]. In the non-conformal cases, we found that the defect entanglement entropy is proportional to the free energy of the theory where it is embedded, but it remains to be seen if this can be written in terms of intrinsic defect quantities.
This work provides a first step toward holographically understanding supersymmetric non-conformal defects in non-conformal theories, a subject which, to the best of our knowledge, remains unexplored.
6.1 On the non-conformal dEE
As mentioned before, maximally supersymmetric Yang-Mills shows a generalized conformal structure [33], in dimensions different from three, if the gauge coupling is allowed to transform under Weyl rescalings. It is therefore natural to expect (5.26) and (5.29) to be written in a form similar to (5.22), that is, in terms of quantities intrinsic to the defect.
Supersymmetric 3D Defect in 5D SYM
As explained before, the connection between 5D maximally supersymmetric Yang-Mills and the 6D $\mathcal{N}=(2,0)$ theory, or equivalently, between the D4-brane in Type IIA and the M5-brane in 11D supergravity, can be used to compute 5D quantities in terms of 6D ones by means of [49].
Codimension-2 conformal monodromy defects in the 6D $\mathcal{N}=(2,0)$ were studied in [24]. There, it was found that the defect entanglement entropy can be written as [^18] This formula is known to hold for a broader class of conformal 4 dimensional defects [8]. Look also at [30] for related works.
$c^{\text{renorm}}_{\text{hol}}=-(4{\cal{A}}^{(4)}-\pi^{2}nh^{(4)}_{D})$, (6.1)
where ${\cal{A}}^{(4)}$ and $h^{(4)}_{D}$ are respectively the Weyl anomaly and the conformal weight of the defect and are explicitly given in [24].
Reducing the background configuration along a direction parallel to the defect leads to the solution in Section 3.4. This reduction fits in the discussion of [49], where it is discussed that the would be lower-dimensional anomaly is given in terms of the higher dimensional one as $\mathcal{A}^{(5)}=e^{\Phi_{0}}(2\pi R)\mathcal{A}^{(6)}$, where $R$ is the size of the S1 in which the theory is compactified, and $e^{\Phi_{0}}$ is the asymptotic value of the lower dimensional dilaton.
If the structure above also exists for the defect anomaly, it is natural to expect the defect entanglement entropy of the non-conformal monodromy defect on 5D maximally supersymmetric Yang-Mills to be
$c_{\text{hol}}^{\text{renorm}}\sim-(4\,\mathcal{I}^{(3)}-\pi^{2}n\,h_{D}^{(3)}),$ (6.2)
where $\mathcal{I}^{(3)}$ is the free energy on the 3-dimensional defect, and $h_{D}^{(3)}$ is the conformal weight of the defect under the generalized conformal structure of maximally supersymmetric Yang-Mills.
Although we aim to report on this later, we note that if we identify
$\begin{split}\mathcal{I}^{(3)}&=\kappa_{d}\frac{5^{4}\pi}{3^{2}2^{15}n^{3}}\left(1+6n+25n^{2}-g^{2}\mu_{F}^{2}\right)\left(1+6n-7n^{2}-g^{2}\mu_{F}^{2}\right)c_{hol}^{(4)},\\ h_{D}^{(3)}&=\kappa_{d}\frac{5^{3}}{3^{2}2^{11}\pi n^{4}}((1+3n)^{2}-g^{2}\mu_{F}^{2})(1-n(3n-2)-g^{2}\mu_{F}^{2})c_{hol}^{(4)},\end{split}$ (6.3)
then (6.2) is satisfied. The reason for providing this ansatz is that the dependence of ${\cal{I}}^{(3)}$ and $h_{D}^{(3)}$ on the defect data $(\mu_{F},n)$ is exactly the same as the one for ${\cal{A}}^{(4)}$ and $h^{(4)}_{D}$ for the defect in the 6D $\mathcal{N}=(2,0)$ theory, given in [24]. We expect this to be the case due to [49]. Moreover, this is also consistent with the fact that for all codimension-2 conformal monodromy defects in maximally supersymmetric theories ([24, 25] and (5.22)) the anomaly/free energy and the conformal weight of the defects can be written as
$\mathcal{A},\mathcal{I}\,\sim\,(y^{\chi}_{\text{core}}-1)c_{\text{hol}}^{(p)},\qquad\qquad h_{D}\,\sim\,\mathcal{C}_{\alpha}c_{\text{hol}}^{(p)}\,\sim\,\sum_{I}q_{I}c_{\text{hol}}^{(p)},$ (6.4)
where $\chi$ is a number that depends on the dimension. We expect this structure to be preserved for codimension-2 monodromy defects in maximally supersymmetric Yang-Mills.
Supersymmetric Line Defect
[…]
It is important to remark that this process does not fit into the description of [49]: at the level of the lower dimensional gauged supergravity, both are 4-dimensional, the difference being the field content and the gauge group. Still, it would be interesting if a relation similar to (5.22) and (6.1) holds to the codimension-2 defect in the D2-brane theory, that is
$c_{\text{hol}}^{\text{renorm}}\sim\mathcal{I}^{(1)}-2\pi nh_{D}^{(1)},$ (6.5)
with $\mathcal{I}^{(1)}$ and $h_{D}^{(1)}$ the free energy and the generalized conformal weight of the defect. This relation has been proven to hold for the codimension-2 monodromy defect in ABJM theory [24], so we are encouraged to think that for the same type of defect in the D2-brane theory (which is also a maximally supersymmetric 3-dimensional theory) this relation should also hold, where the conformal symmetry of the ABJM theory is replaced by a generalized conformal structure.
Following the same structure as (6.4), we propose
${\cal{I}}^{(1)}$ $=\kappa_{d}\frac{49\pi}{96n}\left(16n^{2}-\Delta_{1}^{2/3}\Delta_{2}^{2/3}\Delta_{3}^{2/3}\right)c^{(2)}_{\text{hol}}$,
$h_{D}^{(1)}$ $=\kappa_{d}\frac{35}{128n^{2}}\frac{1}{\Delta_{1}^{1/3}\Delta_{2}^{1/3}\Delta_{3}^{1/3}}\big(8(n+1)n^{2}-(g\mu_{F_{1}}-g\mu_{F_{2}}-2)(2g\mu_{F_{1}}+g\mu_{F_{2}}+2)(g\mu_{F_{1}}+2g\mu_{F_{2}}-2)$
$+2n\left(g^{2}\mu_{F_{1}}^{2}+g^{2}\mu_{F_{1}}\mu_{F_{2}}+g^{2}\mu_{F_{2}}^{2}-4\right)\big)c^{(2)}_{\text{hol}}.$ (6.6)
We hope to report on these proposals in the future. More precisely, using the techniques of [33, 20, 21], we hope to explicitly compute these quantities and to check if (6.2) and (6.5) hold not only for monodromy defects but also for other types of defects [^19] Since this is the case for 2d and 4d conformal defects [51, 52, 8]..
6.2 Outlook
There are other directions we plan to address in the future.
  • So far, we have only dealt with supersymmetric monodromy defects. It would be interesting to check how the defect entanglement entropy or the relations proposed in Section 6.1 are modified by the absence of supersymmetry, which can be achieved by taking $\nu\neq 0$.
  • In [33], it was argued that the generalized conformal structure leads to a modification of the classical Ward identities. It is also known [20, 14] that conformal defects also modify the Ward identities. It would be interesting to (holographically) study the interplay between the contribution of the non-conformal defects to the Ward identities generalized conformal structure.
  • Moreover, it would be interesting to apply the techniques of [33, 53] to compute observables in the non-conformal cases. Moreover, precision holography computations might shed light on the proposed relations (6.3)-(6.5), and the physical meaning of $h_{D}$.
  • The renormalization scheme presented here can be easily extended to the holographic duals to codimension-$n$ monodromy defects in maximally supersymmetric SU($N$) SYM, since in dual frame the metric must take the form
    $\mathrm{d}s^{2}_{\text{dual}}=G(y)\left(\rho^{2}\mathrm{d}x^{2}_{1,p-n}+\frac{\mathrm{d}\rho^{2}}{\rho^{2}}+\frac{1}{P(y)}\mathrm{d}y^{2}+\frac{P(y)}{H(y)}\mathrm{d}s^{2}(\text{S}^{n-1})\right),$ (6.7)
    and asymptote to $\text{AdS}_{p+2}$ written as a foliation of $\text{AdS}_{p+2-n}\times\text{S}^{n-1}$ over an interval. At the level of the $(p+2)$-dimensional gauged supergravity, these solutions contain a (n1)(n-1)-form potential $A_{n-1}\sim c(y)\text{vol}(\text{S}^{n-1})$. This is interpreted as a background form for a (n2n-2)-form symmetry of the dual theory, which in Weyl frame has a non-trivial holonomy on $\text{S}^{n-1}$. Some solutions of this form have already been studied in [54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64].
    The exact same prescription of Section 5.1 works for defects of this form, which allows for a direct computation of the entanglement entropy on the effective theory on these defects.
  • To further study the D5-brane solution presented in Section 4. As argued above, this solution can be interpreted as dual to 6D maximally supersymmetric SU($N$) SYM on $\text{S}^{1}$, similar to [46, 47]. It would be interesting to understand if those two solutions are connected or if they describe two different physical systems.
  • In [34] a solution of 8D U(1) gauged supergravity, corresponding to a D6-brane wrapping a spindle is also considered. We can not use the procedure described in Section 3 to interpret that configuration as a defect, since for the D6-brane the change of coordinates (2.4) and (2.13) sets the UV region at $\rho=0$, so the boundary conditions of Section 3 do not apply here. It would be interesting to construct a solution dual to a codimension-2 defect on the D6-brane theory.
We hope to address these and other possibly related questions in the future.
Acknowledgements
We are thankful to Ignacio Carreño Bolla, Iñaki García Etxebarria, Jerome Gauntlett, Fridrik Gautason, Adolfo Guarino, Yolanda Lozano, Francesco Mignosa, Carlos Nunez, Diego Rodriguez-Gomez, Christopher Rosen, Kostas Skenderis, Jesse Van Muiden, Daniel Waldram and Itamar Yaakov for useful discussion and valuable comments. The work of A.C. is supported by the Severo Ochoa fellowship PA-23-BP22-019. The work of R.S. is supported by the Ramón y Cajal fellowship RYC2021-033794-I. The authors acknowledge support from grants from the Spanish government MCIU-22-PID2021-123021NB-I00, MCIU-25-PID2024-161500NB-I00 and principality of Asturias SV-PA-21-AYUD/2021/52177. AC acknowledges the INGENIUM Alliance of European Universities for giving the opportunity to spend training time at the University of Crete. AC thanks Crete Center for Theoretical Physics and Imperial College London for the kind hospitality while some parts of this work were being completed.
Appendix A Supergravity Conventions
This appendix contains the lagrangians of the different supergravities used throughout this paper. In the lower dimensional cases, we also include the uplift to Type II supergravity.
A.1 Type II Supergravity
We start by reviewing our conventions for Type II supergravity, for which we follow the democratic conventions of [65]. The bosonic lagrangian consists of the NS-NS sector, which contains the metric, dilaton $\Phi$ and NS 3-form $H=dB$, and the RR sector, which contains even(odd) forms in Type II A(B). These expressed using a polyform $F_{\pm}$ defined as
$F_{\pm}=\begin{cases}F_{0}+F_{2}+F_{4}+F_{6}+F_{8}+F_{10}&\text{IIA}\\ F_{1}+F_{3}+F_{5}+F_{7}+F_{9}&\text{IIB},\end{cases}$ (A.1)
so that the upper/lower signs correspond to type IIA/IIB respectively.
In order to have the correct amount of degrees of freedom, the polyform $F$ must also obey a self duality constraint, which halves its degrees of freedom [^20] In the democratic conventions, the Hodge dual is defined such $\star e^{\underline{M}_{1}...\underline{M}_{k}}=\frac{1}{(d-k)!}\epsilon_{\underline{M}_{k+1}...\underline{M}_{d-k}}^{~~~~~~~~~~~~~\underline{M}_{1}...\underline{M}_{k}}e^{\underline{M}_{k+1}...\underline{M}_{d-k}},$ where the $d=10$ indices $M$ are curved and $\underline{M}$ are flat.
$F=\star\lambda(F),$ (A.2)
where $\lambda(C_{k})=(-1)^{[\frac{k}{2}]}C_{k}$.
With these considerations, the equations of motion of type II supergravity, together with the Bianchi identities for the fluxes can be written as
$\mathrm{d}_{H}F_{\pm}=0,~~~~\mathrm{d}H=0,~~~~\mathrm{d}(e^{-2\Phi}\star H)-\frac{1}{2}(F_{\pm},F_{\pm})_{8}=0,$
$2R-H^{2}-8e^{\Phi}(\nabla)^{2}e^{-\Phi}=0,~~~R_{AB}+2\nabla_{A}\nabla_{B}\Phi-\frac{1}{2}H^{2}_{AB}-\frac{e^{\Phi}}{4}(F_{\pm})^{2}_{AB}=0,$ (A.3)
where $(F_{\pm},F_{\pm})_{8}$ is the 8-form part of $F_{\pm}\wedge\lambda(F_{\pm})$, $d_{H}=d+H\wedge$. We have defined for a $k$-form $X$
$(X_{k})_{M}:=\iota_{dx^{M}}X_{k},~~~~X_{k}^{2}:=\sum_{k}\frac{1}{k!}(X_{k})_{M_{1}\dots M_{k}}(X_{k})^{M_{1}\dots M_{k}},$ (A.4)
and for a polyform $X$
$X^{2}_{MN}:=\sum_{k}\frac{1}{(k-1)!}(X_{k})_{MM_{1}\dots M_{k-1}}(X_{k})_{N}{}^{M_{1}\dots M_{k-1}}$. (A.5)
Finally, the SUSY variations for the dilatino and gravitino are given by
$\left(\nabla_{M}-\frac{1}{4}H_{M}\right)\epsilon_{1}+\frac{e^{\Phi}}{16}F_{\pm}\Gamma_{M}\epsilon^{2}=0,\quad\left(\nabla_{M}+\frac{1}{4}H_{M}\right)\epsilon_{2}\pm\frac{e^{\Phi}}{16}\lambda(F_{\pm})\Gamma_{M}\epsilon^{1}=0,$ (A.6a)
$\left(\nabla-\frac{1}{4}H-d\Phi\right)\epsilon_{1}=0,\quad\left(\nabla+\frac{1}{4}H-d\Phi\right)\epsilon_{2}=0.$ (A.6b)
Here we have used the Clifford map for the polyforms. Also, the spin covariant derivative is defined as
$\nabla_{M}=\partial_{M}+\frac{1}{4}\omega_{M}{}^{\underline{P}\underline{Q}}\Gamma_{\underline{P}\underline{Q}},$ (A.7)
with $\omega_{M}^{\phantom{M}\underline{P}\underline{Q}}$ the spin connection.
The spinors $\epsilon_{1,2}$ are Majorana-Weyl spinors. Using the chirality matrix, $\hat{\Gamma}=\Gamma^{\underline{0}...\underline{9}}$, $\epsilon_{1,2}$ satisfy the chirality conditions
$\hat{\Gamma}\epsilon_{1}=\epsilon_{1},~~~~\hat{\Gamma}\epsilon_{2}=\mp\epsilon_{2},$ (A.8)
where, as before, the upper/lower sign corresponds to Type II A/B respectively.
A.2 7D Supergravity and its Uplift
As explained in [34], we are interested in an abelian truncation - common to two models: the 7D ISO(4) and SO(4) gaugings- with two gauge fields $A^{(I)}$ and two scalars $\lambda_{I}$, presented in [66] [^21] Look at [67] for a classification of all possible gaugings in 7d maximal supergravity.. The action of the models is
$S=\frac{1}{16\pi G_{N}^{(7)}}\int\mathrm{d}^{7}x\sqrt{-g}\left[R-2\mathcal{V}-5(\partial{{}_{\mu}}\lambda_{+})^{2}-(\partial{{}_{\mu}}\lambda_{-})^{2}-\frac{1}{4}\sum_{I=1}^{2}e^{-4\lambda_{I}}(F^{I})^{2}\right],\\$ (A.9)
where $\lambda_{\pm}=\lambda_{1}\pm\lambda_{2}$, $F^{I}=dA^{I}$ and the scalar potential is
$\mathcal{V}=-2g^{2}e^{2(\lambda_{1}+\lambda_{2})}.\\$ (A.10)
In this model the R-symmetry gauge field is
$A_{R}=A_{1}+A_{2}.\\$ (A.11)
The uplift to Type IIA/IIB Supergravity has been tidied up in [34]. First we introduce some auxiliary quantities as
$U=e^{2\lambda_{1}}\cos^{2}\eta+e^{2\lambda_{2}}\sin^{2}\eta,\qquad V=2\left[e^{4\lambda_{1}}\cos^{2}\eta+e^{4\lambda_{2}}\sin^{2}\eta-U\left(e^{2\lambda_{1}}+e^{2\lambda_{2}}\right)\right].\\$ (A.12)
These combinations help us to write the ten dimensional metric in string frame metric as
$\mathrm{d}s^{2}_{\text{st}}$ $=e^{2(\lambda_{1}+\lambda_{2})}\mathrm{d}s^{2}_{7}+\frac{1}{g^{2}}\left(\mathrm{d}\eta^{2}+\frac{1}{U}\left(e^{2\lambda_{2}}\cos^{2}\eta(\mathrm{d}\phi_{1}-gA^{1})^{2}+e^{2\lambda_{1}}\sin^{2}\eta(\mathrm{d}\phi_{2}-gA^{2})^{2}\right)\right),$
$B$ $=\frac{1}{2g^{2}U}\left[e^{2\lambda_{1}}\cos^{2}\eta\,\mathrm{d}\phi_{1}\wedge(\mathrm{d}\phi_{2}-2gA^{2})-e^{2\lambda_{2}}\sin^{2}\eta\,(\mathrm{d}\phi_{1}-2gA^{1})\wedge\mathrm{d}\phi_{2}\right],$ (A.13)
$e^{2\Phi}$ $=\frac{e^{6(\lambda_{1}+\lambda_{2})}}{U}.$
where $ds^{2}_{7}$ is the metric of the seven dimensional solution. The expression for $B_{2}$ applies in the case $A^{1}\wedge A^{2}=0$, that is satisfied by the solution of interest. If this is not the case, the uplift is given in terms of its Field Strength tensor $H$, for this we refer to [34].
A.3 6D Supergravity and its Uplift
The solutions in Section 3.4 are obtained using a U(1)2 truncation of the maximally supersymmetric 6D SO(5) gauged supergravity [68]. The bosonic content of this truncation is: the metric, two gauge fields $A^{I}$, and three real scalars $\lambda_{I}$ and $\sigma$, where $I=1,2$. The bosonic lagrangian reads [34]
$S=\frac{1}{16G^{(6)}_{N}}\int\mathrm{d}^{6}x\sqrt{-g}\bigg(R-2{\cal{V}}-5(\partial\lambda_{+})^{2}-(\partial\lambda_{-})^{2}-80(\partial\sigma)^{2}-\sum_{I=1}^{2}\frac{1}{2}e^{-4(\lambda_{I}+\sigma)}(F^{I})^{2}\bigg),\\$ (A.14)
where $F^{I}=dA^{I}$, $\lambda_{\pm}=\lambda_{1}\pm\lambda_{2}$ and the scalar potential is given by
${\cal{V}}=-\frac{g^{2}}{4}e^{4\sigma}\left(8e^{2(\lambda_{1}+\lambda_{2})}+4e^{-4\lambda_{1}-2\lambda_{2}}+4e^{-2\lambda_{1}-4\lambda_{2}}-e^{-8\lambda_{1}-8\lambda_{2}}\right).\\$ (A.15)
We introduce the R-symmetry and flavor gauge field in the same conventions as [24]
$A_{R}=-(A^{2}+A^{1})\qquad A_{F}=-(A^{2}-A^{1})\\$ (A.16)
This theory can be uplifted on $\text{S}^{4}$ to 10D. This uplift can be obtained in the following way [^22] In [69] only the uplift of the metric and the dilaton is derived.: first, the 6D theory is uplifted to 7D U(1)2 gauged supergravity [34] using a U(1) direction $\text{S}^{1}_{\varphi}$. Then, the 7D theory can be uplifted to 11D supergravity using [70]. Finally, the reducing the 11D background on $\text{S}^{1}_{\varphi}$ leads the following Type IIA background
[…]
Here $\alpha=0,1,2$. The scalars $X^{\alpha}$ are related to the 6D ones as
$X^{I}=e^{2\lambda_{I}},\qquad X^{0}=(X^{1}X^{2})^{-2}.$ (A.18)
Also, $\hat{\mu}_{\alpha}$ are the embedding coordinates on S4 such that $\sum_{\alpha=0}^{2}\hat{\mu}^{2}_{\alpha}=1$, and
$\tilde{\Delta}=\sum_{\alpha=0}^{2}X_{\alpha}\hat{\mu}_{\alpha}^{2}.$ (A.19)
Finally, $ds^{2}_{6}$ is the 6-dimensional gauged supergravity metric and vol6 is its volume form.
A.4 5D Supergravity and its Uplift
We now present the 5D U(11)3 gauged supergravity model, following the conventions of [44]. The bosonic content of this model is, the metric, two real scalars $(\phi^{1},\phi^{2})$ and three gauge fields $A^{I}$, $I=1,2,3$. The bosonic lagrangian reads
$S=\frac{1}{16\pi G^{5}_{N}}\int\mathrm{d}^{5}x\sqrt{-g}\Big[R-2\mathcal{V}-\frac{1}{2}\sum_{i=1}^{2}\left(\partial\phi^{i}\right)^{2}-\frac{1}{4}\sum_{I=1}^{3}\left(X^{I}\right)^{-2}(F^{I})^{2}\Big]-F^{1}\wedge F^{2}\wedge A^{3},\\$ (A.20)
where $I=1,2,3$, and $F^{I}=\mathrm{d}A^{I}$. We defined three auxiliary scalars that satisfy the constraints $X^{I}>0$ and $X^{1}X^{2}X^{3}=1$, they can be in terms of $\phi_{1}$, $\phi_{2}$ appearing in (3.17) as
$X^{1}=e^{-\frac{\phi_{1}}{\sqrt{6}}-\frac{\phi_{2}}{\sqrt{2}}}\,,\qquad X^{2}=e^{-\frac{\phi_{1}}{\sqrt{6}}+\frac{\phi_{2}}{\sqrt{2}}}\,,\qquad X^{3}=e^{\frac{2\phi_{1}}{\sqrt{6}}}\,.\\$ (A.21)
Using these auxiliary scalar fields, the potential is given by
$\mathcal{V}=-2\sum_{I=1}^{3}\left(X^{I}\right)^{-1}\,.\\$ (A.22)
To define the R-symmetry and flavour symmetry gauge fields, we follow the same conventions as [20]
$A_{R}=A^{1}+A^{2}+A^{3},\qquad A_{F}=A^{1}-A^{2},\qquad A_{F^{\prime}}=\frac{2}{3}\left(A^{1}+A^{2}-2A^{3}\right).\\$ (A.23)
The uplift of this theory to Type IIB can be found in [70]
$\begin{split}\mathrm{d}s^{2}_{\text{st}}&=\Delta^{1/2}\mathrm{d}s^{2}_{5}+\Delta^{-1/2}\sum_{I=1}^{3}(X^{I})^{-1}\left[\mathrm{d}\hat{\mu}_{I}^{2}+\hat{\mu}_{I}^{2}(\mathrm{d}\phi_{I}+A^{I})^{2}\right],\\ F_{5}&=(1+\star_{10})\Big\{2\sum_{I=1}^{3}\Big[(X^{I})^{2}\hat{\mu}_{I}^{2}-\Delta X^{I}]\text{vol}_{5}+\frac{1}{2}\sum_{I=1}^{3}(X^{I})^{-1}\star_{5}\mathrm{d}X^{I}\wedge\mathrm{d}(\hat{\mu}_{I}^{2})\\ &+\frac{1}{2}\sum_{I=1}^{3}(X^{I})^{-2}\mathrm{d}(\hat{\mu}_{I}^{2})\wedge(\mathrm{d}\phi_{I}+A^{I})\wedge\star_{5}F^{I}\Big\}.\end{split}$ (A.24)
$\hat{\mu}_{I}$ are the embedding coordinates on S5 such that $\sum_{I=1}^{3}\hat{\mu}^{2}_{I}=1$ and
$\Delta=\sum_{I=1}^{3}\hat{\mu}_{I}^{2}X^{I}.$ (A.25)
while $ds^{2}_{5}$ and vol5 are correspond to the 5-dimensional gauged supergravity solution.
A.5 4D Supergravity and its Uplift
We now present the conventions for the solution in Section 3.2 that can be see as the $U(1)^{3}$ truncation of the 4d ISO(7) gauged theory, which can be obtained as a deformation of the maximal SO(8) gauging [40]. The bosonic fields are the metric, three gauge fields $A^{I}$ and four real scalars $\phi_{i}$. The bosonic action is [34]
$S=\frac{1}{16\pi G^{(4)}_{N}}\int\mathrm{d}^{4}x\sqrt{-g}\left[R-2\mathcal{V}-\frac{1}{2}\sum_{i=0}^{3}(\partial\phi_{i})^{2}-\frac{1}{4}\sum_{I=1}^{3}(X^{I})^{-2}(F^{I}_{\mu\nu})^{2}\right]\,$, (A.26)
where we defined the following combinations of scalars
$X^{0}=\frac{1}{2}e^{\frac{1}{2}(\phi_{1}+\phi_{2}+\phi_{3})-\phi_{0}}\qquad X^{1}=e^{\frac{1}{2}(\phi_{1}-\phi_{2}-\phi_{3})},\qquad X^{2}=e^{\frac{1}{2}(-\phi_{1}+\phi_{2}-\phi_{3})},\qquad X^{3}=e^{\frac{1}{2}(-\phi_{1}-\phi_{2}+\phi_{3})},\\$ (A.27)
Using these auxiliary scalar fields, the scalar potential is given by
$\mathcal{V}=-\frac{g^{2}}{2}\left[\sum_{0\leq i<j\leq 3}X^{i}X^{j}-\frac{1}{2}(X^{0})^{2}\right].$ (A.28)
The three gauge fields $A^{I}$ are associated with the $U(1)^{3}\in ISO(7)$ truncation. We can define the R-symmetry and flavour symmetry gauge fields to be
$A_{R}=A^{1}+A^{2}+A^{3},\qquad A_{F_{1}}=A^{1}-A^{2},\qquad A_{F_{2}}=A^{2}-A^{3}.$ (A.29)
The uplift was constructed in [71, 72, 73]. Some details are given in [34].